# Push and Pull: A Framework for Measuring Attentional Agency

### v1

@@ Abstract.
We propose a framework for measuring attentional agency—the ability to allocate one’s attention according to personal desires, goals, and intentions—on digital platforms. Platforms extend people’s limited powers of attention by extrapolating their preferences to large collections of previously unconsidered informational objects. However, platforms typically also allow people to influence one another’s attention. We introduce a formal framework for measuring how much a given platform empowers people to both pull information into their own attentional field and push information into the attentional fields of others. We also use these definitions to shed light on the implications of generative foundation models, which enable users to bypass the implicit “attentional bargain” that underlies embedded advertising and other methods for capturing economic value from informational goods. We conclude with a set of policy strategies that can be used to understand and reshape the distribution of attentional agency online.
@@ 1. Introduction
“The summation of human experience is being expanded at a prodigious rate, and the means we use for threading through the consequent maze to the momentarily important item is the same as was used in the days of square-rigged ships… Consider a future device… in which an individual stores all his books, records, and communications, and which is mechanized so that it may be consulted with exceeding speed and flexibility.”
@@
“In an information-rich world, the wealth of information means a dearth of something else: a scarcity of whatever it is that information consumes. What information consumes is rather obvious: it consumes the attention of its recipients. Hence a wealth of information creates a poverty of attention and the need to allocate that attention efficiently…”
—Herbert Simon, Designing Organizations for an Information-Rich World, 1971
@@
The original promise of the web was to develop a vast interconnected network of information that could, in the words of Tim Burners-Lee, “grow and evolve with the organization and the projects it describes” (Berners-Lee, 1989). Early proponents of expanding the web “world wide” via the internet presumed that knowledge was the primary bottleneck constraining technological ingenuity, economic development, and the spread of democratic values. Collect together and make accessible all the world’s informational resources, the thinking went, and there would be no limit to what humanity could accomplish.
@@
There are a few fundamental challenges with implementing this vision. First, as Herbert Simon (Simon, 1971) so eloquently pointed out, the limitations imposed by human attention constrain the total amount of information that people can actually make use of at any given moment, meaning that the methods by which information is organized, sorted, indexed, and navigated play a central role in determining its practical utility. Indeed, the limitations imposed by human attention mean that adding new information to the internet without concomitant advances in search can actually reduce its overall functionality. Past a certain scale, attentional limitations necessitate curation and, inevitably, grant influence to whoever controls the process by which information is searched. We call this the internet’s attention problem.
The second major complication with the kind of informational web original laid out by Tim Berners-Lee (Berners-Lee, 1989) is that it constitutes what economists call a public good: a resource that is both non-excludable (anyone with access to the internet can benefit from the public information it contains) and non-rivalrous (one person “consuming” information does not reduce its availability to be “consumed” by another person). Anyone who invests personal toil and trouble into creating and maintaining a high-quality content benefits all of humanity, but only receives an economic benefit if they find some way to monetize their contribution—web traffic, on its own, does not pay the bills. In some cases, users will provide content without extrinsic incentives (to satisfy a variety of intrinsic motivations (Redi et al., 2020)). In general, however, public goods tend to be under-provided relative to their socially optimal levels “at a price of zero.” This problem applies both to the creation of valuable content and the provision of technologies that enable one to sort and filter that content. We call this the internet’s incentive problem.
@@
In theory, the internet is a neutral suite of protocols that enable people to freely transmit information between computers on demand. In practice, however, the internet is defined by a collection of platforms that have been built on top of these protocols and that mediate which information flows actually take place. A key insight motivating the present work is that many prominent platform categories (e.g., search engines, content libraries, social media sites, matching platforms, online marketplaces, and, increasingly, consumer AI systems) create value by helping people overcome the attention and incentive problems outlined above. Not only this, but many platforms accomplish this a very similar way—namely, by inferring people’s wants and desires, then extrapolating those preferences to intractably large sets of unconsidered informational objects (e.g., websites, products, people, content, messages, stories, and posts) (Resnick and Varian, 1997; Shapiro and Varian, 1999). To echo Vannevar Bush, platforms have become the “means we use for threading through the consequent maze to the momentarily important item” (Bush et al., 1945).
@@
The choices platforms make about how to infer and extrapolate people’s preferences play a pivotal role in determining who has attentional agency on the internet and, increasingly, the world at large. Informally, we define attentional agency as the ability to distribute attention according to one’s own desires, goals, and intentions. In what follows, we introduce a simple mathematical framework that formalizes and operationalizes this concept. Our framework: (1) clarifies relevant aspects of how various institutions currently solve the attention and incentive problems on the internet; (2) furnishes operational measures of attentional agency that can be used to identify disparities in informational agency across distinct groups of users for any given platform, and (3) lends insight into the implications of generative foundation models and other advances in artificial intelligence.
@@
The fact that many prominent platforms are oriented around solving the same underlying preference inference and extrapolation problem enables us to formalize attentional agency using a common mathematical framework—a generalization of the “discounted cumulative gain” formula that defines the canonical ranking problem and is used in practice as a training objective (Järvelin and Kekäläinen, 2002). The general nature of this approach enables us to abstract away from many idiosyncrasies that distinguish platforms (e.g., whether they sort websites, videos, or restaurants) and cut through to essential questions of equity and access on the modern internet: who has attentional agency on a platform, and why?
Our analysis begins with the observation that many information flows on the internet take the form of a stylized interaction between three parties:
@@
    An advocate seeking to influence what enters the agent’s limited attentional field. Prototypical examples include an advertiser, influencer, or government agency issuing a public-service announcement.
@@
    A platform, which seeks to provide value to both agents (by inferring and extrapolating their preferences to large collections of mostly unconsidered informational objects) and advocates (by allowing them to influence what agents ultimately attend to). Prototypical examples include a search engine, content library, online marketplace, social media platform, and matching platform.
On some platforms, agents and advocates are non-overlapping groups of users (e.g., job-seekers versus employers on a digital marketplace); on others, however, they are the same users acting in different capacities, either simultaneously or at different points in time (e.g., users of a dating app who both view profiles and want their own profile to be widely viewed).
When an agent and advocate have sharply conflicting priorities, a platform cannot deliver maximal value to both simultaneously. An influencer trying to “make a name for themselves,” for example, might want their work to be seen by as many people as possible, while users only want to see content that is narrowly related to their personal interests. We define two separate measures that capture how much attentional value a platform delivers to both agents and advocates.
@@ 1. Pull:
    The amount of attentional value a platform delivers an agent as a fraction of what it is technologically prepared to deliver them given its current knowledge of agent preferences.
@@ 2. Push:
    The amount of attentional value a platform delivers an advocate as a fraction of what it is technologically prepared to deliver them given its current knowledge of agent preferences.
@@
Push and pull are not necessarily zero sum. Indeed, many forms of social interactions allow people to unilaterally enter into the attentional fields of others (e.g., the practice of sending a “cold-email,” which can have net positive value for the recipient). In practice, however, increasing push often does reduce pull on digital platforms: limited attention means that pushing information into an agent’s attentional field tends to exclude other things they would have preferred to see more. As we elaborate in Section 3.2, the relative alignment of agent and advocate preferences and other primitives of a platform environment define the rate at which these two measures trade off against one another.
@@
Our framework defines pull or push at the level of the individual agent or advocate. These estimates can then be used to compare the experience of specific sub-populations as defined by age, race, gender, and other characteristics. Importantly, we define our measures in terms of representations that platforms already compute in the process of mediating informational pushes and pulls—they are therefore not only operationalizable, but in a meaningful sense already operationalized.
@@
Our framework highlights two separate channels by which generative foundation models and other advances in artificial intelligence impact the nature of attentional agency. First, greater intelligence enables platforms to ascertain progressively more granular representation of people’s preferences and intentions leading up to instances of information retrieval. From the perspective of attentional agency, this is potentially a double-edged sword—more accurate representations of a user’s type not only enables platforms to deliver more targeted informational pulls, but also more targeted informational pushes. Our framework reveals that the ultimate agency implications of these innovations depends on the interaction of a variety of factors, including the alignment of intentions between those seeking to push and pull information and the relative accuracy with which platforms can ascertain informational relevance for each.
@@
Second, generative foundation models excel at inferring the significance of complex digital objects at the level of their atomic informational components (“tokens”) and reconstituting them in novel ways (Bommasani et al., 2021; Bubeck et al., 2023; Yin et al., 2023). Generative foundation models therefore enable platforms to source, extract, and recombine information more flexibly. For example, whereas traditional search engines ranked relevant web pages in their entirety, AI-powered search tools can pull, rank, sort, and recombine information from within a variety of different pages on demand. This poses a fundamental challenge to the existing social, economic, and legal institutions that have emerged to regulate the capture and distribution of value on the internet. Most notably, the ability of generative foundation models to surgically extract information from its original delivery format—thus leaving embedded advertising, branding, and other promotional material behind—undermines many existing methods that platforms and content providers rely on to monetize their services. This is especially disruptive of the advertising model, which relies on the fact that people apprehend traditional informational bundles (e.g., web pages) in their entirety. Stripping away context can also undermines established norms and expectations about how information should be interpreted, modified, and shared (Jain et al., 2023).
@@
The rest of the paper is organized as follows. In Section 2, we first review related literature. Section 3 then presents our framework and key definitions. Section 4.1 highlights the broad applicability of our framework by detailing how five major categories of current internet platforms—search, content, social media, matching, and marketplaces—map onto the primitives of our model. We then turn to the question of how generative foundation models will impact attentional agency on the internet in Section 4.2. Section 5 presents some of the immediate policy implications of the concepts and formal definitions we propose and highlights open research questions.
@@ 2. Related Work
2.1. Algorithmic Harms and Fairness
The potential harms of digital platforms have become a central topic in modern computer science (see Shelby et al. for an overview and taxonomy (Shelby et al., 2023)). Similar to our goal of operationalizing attentional agency on digital platforms, much of this work seek to translate notions such as broad beneficence, equity, and fairness into new technological domains to ensure that they are upheld (Kazim and Koshiyama, 2021).
For example, Benthall and Shekman (Benthall and Shekman, 2023) propose to extend legal fiduciary duties to the design of artificial intelligence systems by requiring that they both accurately learn people’s preferences and respect concepts such as loyalty and care. Bar-Gill et al. argue that algorithms can accelerate the capacity of firms and other sophisticated actors to exploit identifiable groups (Bar-Gill et al., 2023). Similarly, Rakova et al. propose to introduce “meaningful meaningful consent and contestability” into the legal and social relationship between online platforms and users.
Our measures of attentional agency support these existing approaches to aligning platforms to user preferences. The general perspective we develop also dominance of information ranking on the modern internet, which highlights the relevance of theoretical work on fair ranking (e.g., (Hafalir et al., 2021; Patro et al., 2022)) for broader issues of equity and access on the internet.
@@ 2.2. Economic, Psychological, and Social Value of Attention
People’s preferences over what they pay attention to arise from a wide variety of psychological, social, and economic considerations. To take but a few examples:
  • Emotional: What we pay attention directly affects our well-being through hedonic states such as curiosity, boredom, flow, mental effort, disgust, outrage, anxiety, excitement, jealousy, and regret (Loewenstein and Wojtowicz, 2023).
  • Cognitive: Paying attention to what others attend to (“joint attention”) is a deep psychological primitive that drives childhood development and adult social interactions (Mundy and Newell, 2007; Frischen et al., 2007). Simultaneously attending to common objects in physical or digital space is a key part of what makes our shared reality shared as such (Tomasello et al., 2005; Chwe, 2013).
  • Social: Receiving positive attention from others is both socially advantageous and psychologically gratifying (Braudy, 1997; Meshi et al., 2013; Lindström et al., 2021). Conversely, being ostracized or publicly shamed are among the most painful experiences one can endure (Zadro et al., 2004; Dickerson et al., 2004). The “fear of missing out” and related emotions, in particular, seem to play a key role in motivating social media use (Bursztyn et al., 2023; Abel et al., 2016).
  • Decisional: Attention is necessary to both absorb and process decision-relevant information (Stigler, 1961).
  • Economic: One’s attention can be rented to firms (the basis of many “white collar” jobs).
  • Educational: Attention is necessary to learn and make other investments in one’s human capital, especially through formal education.
Our framework distills these diverse motivations into to a single utility function and asks: to what degree to platforms maximize people’s attentional priorities given their current information and technology?
@@
By operationalizing attentional agency in terms inferred utility values, our framework inherits both the strengths and weaknesses of revealed preference analysis. On the one hand, this approach enables us to capture a wide variety of motivations in a single measure that can be inferred from people’s behavior using standard techniques (indeed, the very techniques that platforms already use to deliver value to users). On the other hand, our measures leave open the question of how tightly coupled people’s attentional preferences are to their personal well-being or broader notions of social welfare.
@@
Indeed, there are many situations where the two seem to diverge quite sharply: attentional addiction; highly engaging but politically divisive content; people seeking information with the intention of harming themselves or others; etc. Although our framework furnishes an operational measure of attentional agency, the question of how to evaluate its desirability in various circumstances is nuanced and requires careful consideration. Some of these concerns are only becoming more acute as AI search and other innovations enable platforms to deliver narrowly targeted persuasive messaging (Durmus et al., 2024; Salvi et al., 2024) and other novel forms of harmful content to users—an issue we return to in Section 5.2.
2.3. Information Retrieval
Our approach is motivated by the observation that information retrieval plays a central role in modern internet platforms (see Table 1). Solutions to this problem follow a common approach: (1) assign user-specific “relevance” scores to various pieces of informational content, (2) use those scores to sort content; (3) arrange items in an interface in such a way as to preferentially guide attention to items with higher rankings. Formal methods for generating and evaluating the quality of ranked content lists, known as “recommender systems,” have long been a topic of study in theoretical and applied computer science (Herlocker et al., 2004).
Approaches to recommendation include measuring the precision and recall of a list of a fixed size (e.g., the “ten blue links” returned by a search engine), considering the mean reciprocal rank (applicable to social media feeds of arbitrary extent), and variants of the discounted cumulative gain, which calculate a weighted average of relevance over ranks. Researchers have also explored methods for understanding how allocation impacts item diversity, bias, and other features of interest (Parapar and Radlinski, 2021; McAuley, 2022). The work of Ref. (Stray et al., 2021), in particular, points out that recommender systems are a special case of the more general value alignment problem and emphasizes the importance of developing operational measures of alignment—the central task of this paper.
Although the exact correspondence between offline ranking metrics and company revenues is typically proprietary, some researchers have sought to describe how offline metrics map to business value, ultimately concluding (as we might expect) that whether a given ranking metric works well is context-dependent (Jannach and Jugovac, 2019). Published work from Netflix has also discussed the role of offline experimentation and the use of offline ranking metrics like NDCG to predict business value (Alvino and Basilico, 2015; Gomez-Uribe and Hunt, 2016).

### v2

@@ Abstract.
We propose a framework for measuring attentional agency, which we define as a user’s ability to allocate attention according to their own desires, goals, and intentions on digital platforms that use statistical learning to prioritize informational content. Such platforms extend people’s limited powers of attention by extrapolating their preferences to large collections of previously unconsidered informational objects. However, platforms typically also allow users to influence the attention of other users in various ways. We introduce a formal framework for measuring how much a given platform empowers each user to both pull information into their own attention and push information into the attention of others. We also use these definitions to clarify the implications of generative foundation models and other recent advances in AI for the structure and efficiency of digital platforms. We conclude with a set of possible strategies for better understanding and reshaping attentional agency online.
Attention, agency, ranking, recommender systems, digital platforms, social media, search, generative AI, human-computer interaction, information retrieval
journalyear: 2025copyright: acmlicensedconference: The 2025 ACM Conference on Fairness, Accountability, and Transparency; June 23–26, 2025; Athens, Greecebooktitle: The 2025 ACM Conference on Fairness, Accountability, and Transparency (FAccT ’25), June 23–26, 2025, Athens, Greecedoi: 10.1145/3715275.3732043isbn: 979-8-4007-1482-5/2025/06ccs: Information systems Retrieval models and rankingccs: Applied computing Law, social and behavioral sciencesccs: Information systems Personalization
@@ 1. Introduction
“The summation of human experience is being expanded at a prodigious rate, and the means we use for threading through the consequent maze to the momentarily important item is the same as was used in the days of square-rigged ships… Consider a future device… in which an individual stores all his books, records, and communications, and which is mechanized so that it may be consulted with exceeding speed and flexibility. It is an enlarged intimate supplement to his memory.”
@@
The original promise of the web was to create an interconnected network of information that could, in the words of Tim Berners-Lee, “grow and evolve with the organization and the projects it describes” (Berners-Lee, 1989). Early advocates of expanding the web “world wide” via the internet argued that limited access to information was a substantial bottleneck constraining both technological innovation and economic development. It seemed to follow logically that making the world’s informational resources accessible online would greatly expand what humanity could accomplish.
In the decades since, the internet has grown precipitously, yet its originally envisioned benefits remain only partially realized. As it turns out, simply increasing the volume of freely available information online has not automatically produced a commensurately more well-informed global populace, in part because attentional constraints sharply limit what people can actually absorb and process (Loewenstein and Wojtowicz, 2023). As Herbert Simon (1971) so presciently pointed out in the early days of the information age, decades before the world wide web was introduced:
“In an information-rich world, the wealth of information means a dearth of something else: a scarcity of whatever it is that information consumes. What information consumes is rather obvious: it consumes the attention of its recipients. Hence a wealth of information creates a poverty of attention and the need to allocate that attention efficiently.”
@@
Thus, attentional limitations mean that the practical utility of the internet depends critically on our methods for organizing and navigating the vast quantities of information it contains. Indeed, the limitations imposed by human attention mean that adding information to the internet without concomitant advances in search technology can actually reduce its overall functionality. Past a certain scale, attentional limitations necessitate curation, which, in turn, inevitably grants influence to whoever controls the curatorial process. We call this the internet’s fundamental attention problem.
@@
In theory, the internet is a neutral suite of protocols that enable people to freely transmit information between computers on demand. In practice, however, a collection of platforms built on top of these core protocols mediate which information flows actually take place. A key insight motivating the present work is that many prominent platform categories (e.g., search engines, content libraries, social media sites, matching platforms, online marketplaces, and, increasingly, consumer AI systems) all create value by helping people overcome the internet’s fundamental attention problem outlined above.
In fact, many platforms accomplish this goal in essentially the same way—namely, by inferring people’s attentional preferences, then algorithmically extrapolating them to intractably large sets of unconsidered informational objects (e.g., websites, products, people, content, messages, stories, and posts; see Table 1) (Resnick and Varian, 1997; Shapiro and Varian, 1999). To echo Vannevar Bush, platforms have become the “means we use for threading through the consequent maze to the momentarily important item” (Bush, 1945).
@@
The algorithms that platforms employ to infer and extrapolate users’ attentional preferences play a pivotal role in shaping who has agency on the internet and, increasingly, the world at large. In this paper, we introduce a formal framework for measuring the attentional agency that a platform accords to each of its users. We define attentional agency as the capacity to allocate attention according to personal desires, goals, and intentions. Our framework offers a mathematical formulation of this concept that: (1) clarifies how various institutions currently solve the internet’s fundamental attention problem, (2) furnishes operational measures of attentional agency that can be used to identify differences across users for any given platform, and (3) lends insight into how generative foundation models (GFMs) and other advances in artificial intelligence will impact the current paradigm for matching information and attention on the internet.
1.1. Summary of our framework
In using the term “agency” throughout this paper, we intentionally maintain a narrow focus on people’s capacity to influence what they and others see on a given digital platform, acknowledging that this leaves out other potentially rich notions of agency (e.g., the specific actions a platform makes available to users). [^1] We do not explicitly model feedback loops by which people’s preferences are changed by their exposure to platforms and ranking systems. This is a critical area of future work, but as we will discuss below we believe even a relatively “static” analysis of attentional agency as described here provides immediate value. The motivation for this approach is that it yields measures that are immediately calculable using information that platforms already gather in the process of delivering services to users, and can therefore be readily implemented by technologists and policymakers. [^2] Our core approach and definitions are more general than this, however, as they can be applied to any utility function over informational objects for agents and advocates.
@@ 1.1. Summary of our framework
The fact that many prominent platforms are oriented around solving the same underlying preference inference and extrapolation problem enables us to formalize attentional agency using a common mathematical framework—a generalization of the “discounted cumulative gain” measure used in ranking systems (Järvelin and Kekäläinen, 2002). The general nature of this approach enables us to abstract away from many idiosyncrasies that distinguish platforms (e.g., whether they sort websites, videos, or restaurants) and address a fundamental question: who has attentional agency? Specifically, our analysis begins with the observation that many information flows on the internet take the form of a stylized interaction between three parties:
@@
    An advocate seeking to influence what enters the agent’s limited attentional field. Prototypical examples include an advertiser, government agency issuing a public-service announcement, or influencer.
@@
    A platform, which seeks to provide value to both agents (by inferring and extrapolating their preferences to large collections of mostly unconsidered informational objects) and advocates (by allowing them to influence what agents ultimately attend to). Prototypical examples include search engines, content libraries, online marketplaces, social media platforms, matching platforms, and, increasingly, consumer-focused generative AI products.
On some platforms, agents and advocates are non-overlapping groups of users (e.g., job-seekers versus employers on a job board); on others, however, they are the same users acting in different capacities, either simultaneously or at different points in time (e.g., users of a dating app who both view profiles and want their own profile to be widely viewed).
As we show in greater detail below, when an agent and advocate have sharply conflicting priorities, a platform often cannot deliver maximal value to both simultaneously. An influencer trying to “make a name for themselves,” for example, might want their work to be seen by as many people as possible, while other users only want to see content that is related to their personal interests. We define two separate measures that capture how much attentional value a platform delivers to both agents and advocates.
@@ 1. Pull:
    The amount of attentional value a platform delivers to an agent as a fraction of what it is technologically prepared to deliver them given its current knowledge of agent preferences.
@@ 2. Push:
    The amount of attentional value a platform delivers to an advocate as a fraction of what it is technologically prepared to deliver them given its current knowledge of agent preferences.
@@
Push and pull are not necessarily zero sum. Indeed, many forms of social interactions allow people to unilaterally enter into the attention of others (e.g., the practice of sending a cold-email, which can have net positive value for the recipient). In practice, however, increasing push often does reduce pull on digital platforms: limited attention means that pushing information into an agent’s attention tends to exclude other objects they would have preferred to see more. As we elaborate in Section 3.2, the relative alignment of agent and advocate preferences and other primitives of a platform environment determine the rate at which these two measures trade off against one another.
@@
Our framework defines pull and push at the level of an individual user for a given platform. These estimates can then easily be aggregated across users or systems. For instance, our push and pull measurements could be used to measure the attentional agency of specific sub-populations as defined by community membership, age, race, or other characteristics. Importantly, we define our measures such that the representations that platforms already compute in the process of mediating informational pushes and pulls can be used to quickly estimate different levels of agency—they are therefore not only operationalizable, but in a meaningful sense already operationalized.
@@
Our framework highlights two separate channels by which GFMs and other advances in AI are impacting the nature of attentional agency online. First, platforms will ascertain progressively more granular representations of people’s preferences and intentions. From the perspective of attentional agency, this is a double-edged sword—more accurate representations of a user’s type not only enable platforms to deliver more targeted informational pulls, but also more targeted pushes.
@@
Second, GFMs excel at inferring the significance of complex digital objects at the level of their atomic informational components (“tokens”) and reconstituting them in novel ways (Bommasani et al., 2021; Bubeck et al., 2023; Yin et al., 2023). These models therefore enable platforms to extract and recombine information more flexibly. Whereas traditional search engines rank relevant web pages in their entirety, AI-powered search tools now recombine information from a variety of different pages into new informational objects on demand. This poses a fundamental challenge to the existing social, economic, and legal institutions that have emerged to generate and capture value on the internet. Most notably, the ability of GFMs to surgically extract information from its original delivery format—thus leaving embedded advertising, branding, and other promotional material behind—undermines many existing methods that platforms and content providers rely on to monetize their services. This is especially disruptive of the digital advertising model, which relies on the fact that people apprehend traditional informational bundles (e.g., web pages) as an integral whole. Stripping away context can also undermine established norms and expectations about how information should be interpreted, modified, and shared (Jain et al., 2023).
@@
The rest of the paper is organized as follows. In the rest of Section 1, we provide a high-level summary of our framework, with additional motivation and an early coverage our limitations. In Section 2, we review related literature. Section 3 then presents our framework more formally and develops a mathematical analysis of our definitions. Section 4 highlights the broad applicability of our framework to five major categories of current internet platforms—search, content, social media, matching, and marketplaces. Section 5 presents some of the immediate policy implications of the concepts and formal definitions we propose and highlights open research questions.
@@ 2. Related Work
Our work builds on ideas from a variety of literatures in computer science, economics, and behavioral science. Below, we summarize key prior work, focusing first on proposals to better align digital platforms. Next, we discuss how our framework builds on prior literature evaluating the efficacy and fairness of ranking systems. Finally, we summarize work from psychology and economics on the analysis of individual preferences, which form a key element of our approach.
2.1. Aligning Digital Platforms
The challenge of aligning digital platforms has emerged as a key concern in modern computer science, prompting a variety of authors to propose methods for operationalizing potential social objectives, such as beneficence, equity, and fairness (Kazim and Koshiyama, 2021; Shelby et al., 2023). The growing capabilities of GFMs have, moreover, introduced new forms of misalignment and new risks associated with failing to achieve alignment. A variety of potential solutions have been proposed in the literature, such as subjecting AI systems to a fiduciary requirement that would require them to both accurately learn people’s preferences and uphold general principles, such as loyalty and care (Benthall and Shekman, 2023). Other related proposals include granting users new avenues of contestability (Rakova et al., 2023) and implementing fair ranking in domains like education (Hafalir et al., 2021; Bar-Gill et al., 2023).
Conceptually, our article responds to a call originally made by Stray et al. (2021), who argue that recommender systems represent a specific instance of this broader value alignment problem. They specifically advocate for the creation of operational measures of alignment—developing such a measure is the central aim of this paper.
2.2. Evaluation of Ranking Systems
Our approach centers on a key feature of existing ranking and recommendation systems (Herlocker et al., 2004), namely the user-specific “relevance” scores that platforms use to match user attention and informational content. Existing approaches to evaluating recommender systems include measuring the precision and recall of a list of a fixed size (e.g., the “ten blue links” returned by a search engine), considering the mean reciprocal rank (applicable to social media feeds of arbitrary extent), or calculating variants of the discounted cumulative gain, which calculate a weighted average of relevance over ranks.
Our work intersects with prior research aimed at understanding how the design of recommender systems influences outcomes of interest, such as item diversity, bias, and user satisfaction (Parapar and Radlinski, 2021; McAuley, 2022). Given that the business value of ranking is typically proprietary and somewhat difficult to measure, estimates vary widely (Jannach and Jugovac, 2019). Published work from Netflix suggests that the benefits of ranking-based personalization can be quite large in practice, at least in some domains (Alvino and Basilico, 2015; Gomez-Uribe and Hunt, 2016).
Our measure of attentional agency advances the literature on ranking evaluation by directly mapping this construct to relevance scores using a calculation that can be easily performed using existing infrastructure and institutional practices.
2.3. Fair Ranking
A large body of literature in machine learning has highlighted techniques to make ranking systems more “fair,” as interpreted in a variety of ways (Dwork et al., 2012; Hardt et al., 2016; Kleinberg et al., 2016; Diakopoulos, 2016; Wang et al., 2023).
Patro et al. (Patro et al., 2022) critically review a large body of work focused on fairness for ranking, recommendation, and retrieval systems. They highlight the inherent challenge of defining fairness in the face of a “gap between high ranking placements and true provider utility, spillovers and compounding effects over time, induced strategic incentives, and the effect of statistical uncertainty” (pg. 1929). Relatedly, Chen et al. (Chen et al., 2023) provide a synthesis of how “bias” has been studied in the context of recommendation systems, arguing that group unfairness is but one of seven types of bias.
Recent literature on fair ranking has underscored the difficulty of promoting fairness through a single definition, especially given that perception of fairness can vary widely from person to person (Alkhathlan et al., 2024). Techniques for incorporating fairness into ranking also tend to face a trade-off between fairness and utility (albeit techniques to ameliorate such trade-offs have also been developed (Dinh et al., 2024)). One goal of our proposal is to draw the fairness community’s attention to attention—that is, to highlight attentional agency as a potentially important factor in people’s appraisals and experiences of fairness on digital platforms.
@@ 2.4. Economic, Psychological, and Social Value of Attention
People’s preferences over what they pay attention to arise from a wide variety of emotional (Loewenstein and Wojtowicz, 2023), cognitive (Tomasello et al., 2005), social (Braudy, 1997), decisional (Stigler, 1961), and other considerations. Our framework distills these diverse motivations into a single utility function and asks: to what degree do platforms empower people to satisfy their own attentional priorities?
@@
By operationalizing attentional agency in terms of inferred utility values, our framework inherits both the advantages and limitations of revealed preference analysis. We adopt this approach because it enables us to capture a wide variety of motivations in a single measure that can be inferred from people’s behavior using standard techniques (indeed, the very techniques that platforms already use to deliver value to users).
@@
Revealed preference measures do, however, leave open the question of how tightly coupled people’s attentional preferences are to their personal well-being or broader notions of social welfare (Kleinberg et al., 2024). Indeed, there are many situations where the two seem to diverge quite sharply, such as video game addiction or “doom scrolling” highly engaging but emotionally deleterious content. Although our framework furnishes an operational measure of attentional agency, the question of how to evaluate its desirability in various circumstances is nuanced and requires careful consideration. Some of these concerns are only becoming more acute as AI search and other innovations enable platforms to deliver narrowly targeted persuasive messaging (Durmus et al., 2024; Salvi et al., 2024) and other novel forms of harmful content to users.